OpenStudy (anonymous):

PLEASE HELP A 7.5 kg block initially at rest is pulled to the right along a horizontal, frictionless surface by a constant, horizontal force of 19.6 N. Find the speed of the block after it has moved 2.4 m. Answer in units of m/s I got 1.77 but it's wrong?

OpenStudy (anonymous):

you first need to work out your acceleration before you can work out your velocity

OpenStudy (anonymous):

U(initial velocity) = 0 ; F = 19.6N; S = 2.4m; M = 7.5kg calculate V = ?

OpenStudy (anonymous):

Actually I just used the equation \[k=.5mv^2\] :)

OpenStudy (anonymous):

F = MA 19.6N = 7.5kg(a) to get the (a) on it own do the following: \[\frac{ 19.6 }{ 7.5 } = \frac{ 7.5 }{ 7.5 }a\] 2.613 = a so a = 2.613 m/s now we can calculate velocity/speed: \[s = \frac{ v ^{2} - u 2 }{ 2a }\] \[2.4m = \frac{ v ^{2} - \left( 0 \right)^{2} }{ 2(2.613) }\] \[2.4m = \frac{ v ^{2} }{ 5.226 }\] now to have \[v ^{2}\] alone you must * both sides by \[\frac{ 1 }{ 5.226 }\] when you canceled out 5.226 on the right side , multiply on the left, which will leave you with : 10.452 = \[10.452 = v ^{2}\] now take the \[\sqrt{10.452} = \sqrt{v ^{2}}\] which will leave you with the answer... \[3.233 = v\]

OpenStudy (anonymous):

did you get the same answer as mine?

OpenStudy (anonymous):

I got 3.54 but pretty close :) thanks

OpenStudy (anonymous):

cool. thanks